OpenStudy (anonymous):

1) WILL BE GIVING MEDAL AND FAN!!! Use the quadratic formula to solve 2y2 – 10y + 8 = 0. a. {–4, –1} b. {4, –1} c. {–4, 1} d. {4, 1}

OpenStudy (kidrah69):

First list out your a b c values

OpenStudy (kidrah69):

what are they?

OpenStudy (anonymous):

ok..I know how but....?

OpenStudy (anonymous):

ill do them rn

OpenStudy (zale101):

Factor out the common factor first.

OpenStudy (zale101):

Then work your way on factoring the rest

OpenStudy (anonymous):

a. 2 b. 10 c. 8

OpenStudy (kidrah69):

it says to use the quadratic formula...not factor

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

ohhh then im not sure sat all

OpenStudy (zale101):

Oh, i just noticed that xD

OpenStudy (kidrah69):

b is wrong

OpenStudy (anonymous):

would it be A?

OpenStudy (kidrah69):

No the value

OpenStudy (zale101):

@brae , do you know what the quadratic formula look like?

OpenStudy (kidrah69):

\[\frac{ x = -b ± \sqrt(b^2 - 4ac) }{ 2a } \]

OpenStudy (anonymous):

ummm...wait one sec, I have it in my notes

OpenStudy (kidrah69):

lol i posted it XD

OpenStudy (anonymous):

ohh thats right!!

OpenStudy (zale101):

kidrah, the x shouldn't be in the numerator :X

OpenStudy (kidrah69):

my bad :P doesnt make a diffrence >_<

OpenStudy (zale101):

So @brae , what does the a, b, and c indicate in the quadratic formula, kidrah listed ?

OpenStudy (anonymous):

isnt it this? a. 2 b. 10 c. 8

OpenStudy (zale101):

Do you have any idea for where constants came from?

OpenStudy (kidrah69):

B is still wrong

OpenStudy (anonymous):

yes I learned it today

OpenStudy (kidrah69):

(2y^3) (-10y) (+8)

OpenStudy (kidrah69):

its not 10

OpenStudy (anonymous):

ohh is B -10

OpenStudy (kidrah69):

mhm

OpenStudy (anonymous):

yay!

OpenStudy (kidrah69):

\[\frac{ x = -(-10) ± \sqrt((-10)^2 - 4(2)(8) }{ 2(2) } \]

OpenStudy (kidrah69):

now you just have to solve

OpenStudy (zale101):

The standard quadratic equation is ually written like this \(\Large ax^2+bc+c=0\) and there's a proof for this standard equation that converts to the qudratic formula where you solve x, only.

OpenStudy (anonymous):

Is the answer A??

OpenStudy (anonymous):

i think lol

OpenStudy (anonymous):

maybe not ha

OpenStudy (zale101):

But since your variable is y \(ay^2+by+c=0\) \(2y^2 – 10y + 8 = 0\) a=? b-? c=? \(y=\LARGE \frac{-b\pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (zale101):

Plug and chug

OpenStudy (anonymous):

ok...

OpenStudy (zale101):

and show your work :)

OpenStudy (anonymous):

ummm ok i messed up and is really confused lol

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

Show me what you had.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

but its weird because i learned it today in class

OpenStudy (zale101):

Here is a good example of a quadratic equation being plugged into the quadratic formula. http://www.mathplanet.com/images/math/codecogs_d3fa4daa.gif plug and chug, and do some calculator crunches

OpenStudy (anonymous):

ohh ok so the answers there was 5 and -2

OpenStudy (zale101):

This is just an example i pulled off of google, but try to see how they got those two answers.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

does anyone know the answer?

OpenStudy (anonymous):

@Zale101 @kidrah69

OpenStudy (wolf1728):

2y2 – 10y + 8 = 0 a=2 b=-10 c=8 x= [--10 +- sqroot (100 - 4*2*8)]/2*2 x = [10 +- sqroot(36)] /4 x = [10 + 6] / 4 x1= 4 x2 = [10 - 6] / 4 x2 = 4 / 4 x2=1 Answer is d

OpenStudy (anonymous):

can i tag u in other problems that i have to do?

OpenStudy (wolf1728):

okay - as long as it's not too many others How about just one more problem?

OpenStudy (anonymous):

ya thanks!